What is the solubility of M(OH)2 in a 0.202 M solution of M(NO3)2? Ksp = 6.65×10?18 for M(OH)2 Calculate the molar solubility of lead thiocyanate in 0.600 M KSCN.? Ksp=2.00* 10^-5
Answer #1

Both the problems are examples common ion effect.

Let x be the molar solubility of M(OH)2. The dissociation of M(OH)2 can be written as

M(OH)2 (s) <======> M2+ (aq) + 2 OH- (aq)

x                                      x                   2x

Again, the ionization of M(NO3)2 can be written as

M(NO3)2 (aq) ---------> M2+ (aq) + 2 NO3- (aq)

0.202 M                        0.202 M

We do not bother about the concentration of NO3- since it doesn’t play any role in the precipitation reaction. We next set up the ICE chart as

M(OH)2 (aq) <=====> M2+ (aq) + 2 OH- (aq)

initial                         x                                     0.202               0

change                    - x                                      + x                + 2x

equilibrium               -                                   (0.202 + x)         2x

We need to account for the M2+ furnished by the M(NO3)2. The solubility product is

Ksp = (0.202 + x)(2x)2

Next, we make an approximation. Since the value of Ksp = 6.65*10-18 is quite low, we will expect x<<0.202 and hence, (0.202 + x) ≈ 0.202. Therefore,

Ksp = (0.202).(2x)2

===> 6.65*10-18 = 0.202*4x2

===> x2 = 8.230*10-18

===> x = 2.868*10-9

The molar solubility of M(OH)2 in 0.202 M M(NO3)2 is 2.868*10-9 M ≈ 2.9*10-9 M (ans).

We follow the same logic as before. The concentration of SCN- from KSCN is 0.600 M [1:1 dissociation as per the equation KSCN (aq) ----------> K+ (aq) + SCN- (aq)].

The ionization of lead thiocyanate, Pb(SCN)2 follows the equation

Pb(SCN)2 (s) <=====> Pb2+ (aq) + 2 SCN- (aq)

x                                      x                 2x

Ksp = (x).(2x + 0.600)2

Now, we shall again assume (2x + 0.600) ≈ 0.600; therefore,

Ksp = x.(0.600)2

===> 2.0*10-5 = x.(0.36)

===> x = 5.555*10-5

The molar solubility of Pb(SCN)2 in 0.600 M KSCN is 5.555*10-5 M ≈ 5.5*10-5 M (ans).